OpenStudy (anonymous):

Factor the expression completely (show work) 5x^2-19x+12

OpenStudy (anonymous):

What i am not understanding is the factors of 12 do not add up to get -19, that is what i need help with.

OpenStudy (ilovecake):

5 * 12 = 60 two factors of 60 that add to get -19? -4 and -15. so 0 = 5x^2 - 15 x - 4x + 12 do some factoring 0 = 5x(x - 3) - 4 (x - 3) now you see how have a common (x - 3) so you can factor that out. 0 = (x - 3) (5x - 4) setting each one to 0 and solving for x, you get x = 3, -4/5

OpenStudy (ilovecake):

does that help you?

OpenStudy (anonymous):

so we are multiplying the 5x * 12 to get 60 then finding the factors of 60 that add up to -19?

OpenStudy (ilovecake):

Yes.

OpenStudy (anonymous):

okay i need an answer that looks like (x +_)(x__)

OpenStudy (anonymous):

(x-4)(-15) ??

OpenStudy (ilovecake):

yup

OpenStudy (anonymous):

awesome! Thank you!

OpenStudy (ilovecake):

Np!